OpenStudy (anonymous):

You meet a nasty dog in a dark alley. The dog has a bag and the bag has a pebble in it. The pebble is equally likely to be black or white. The dog then adds a white pebble and shakes up the bag. He then asks you to take a pebble from the bag without looking inside it. What are your chances of taking out a white pebble? (Taken from Richard Wiseman's blog). What if the odds of the original pebble being black or white weren't supplied?

OpenStudy (anonymous):

0.75

OpenStudy (anonymous):

Prove it!

OpenStudy (anonymous):

There are two cases possible. First, the bag already had a white ball, in which the probability of picking a white ball is 1 or 100%. The second case is when the bag has a black ball. So the probability of picking a white one is 0.5 (50%). The combined probability is the sum divided by two. (1+0.5)/2 = 0.75

OpenStudy (anonymous):

Nicely done! Much simpler than my proof: P(picking white pebble) = P(initial pebble white) * P(pick white pebble in this case) + P(initial pebble not white) * P(pick white pebble in this case). In shorthand: \(P(W) = P(A) * P(W/A) + P(A’) * P(W/A’)\) P(W/A) is always 1 (picking a white pebble from a bag of two white pebbles), P(W/A’) is 0.5 (picking a white pebble from a bag of one white and one black), P(A’) = 1 – P(A) So we get: \(P(W) = P(A) * 1 + (1 – P(A)) * 0.5;\) \(P(W) = 1/2 (1 + P(A))\) \(P(W) = 1/2 (1+0.5)\) = 0.75

OpenStudy (anonymous):

Your answer is good even when we don't know the probability of the initial ball being white or black.

OpenStudy (anonymous):

Yeah, I must have intentionally left the P(A) in until the end. Did it a few weeks ago, and don't really remember it!